# Distributed and heterogeneous tensor–vector contraction algorithms for high performance computing

### v1

@@ 5.5 Mixed precision effects on performance
Since both dTVC and dHOPM remain memory bounded algorithms, mixed precision can be employed to further increase their throughput. In fact, mixed precision is a very popular technique actively used in deep learning applications [34, 35]. In this work, low precision is used exclusively as a storage format and we promote numbers to high precision, i.e. by doubling their bit representation, just before computing them. Therefore, every arithmetic operation, besides accumulations, is done in high precision. Calculated quantities are then converted back to their original storage precision. It is worth noting that mixed precision also affects communication and hence requires the development of ad-hoc MPI functions that enable high precision scalar and array sums while keeping data transfers in low precision.
We consider three mixed precision formats: single-double, half-single, and brain-single. The half format stands for the IEEE 754-2008 binary16 standard while brain consists of a 16-bit truncation of the binary32 type from the same standard. While it is straightforward to create a fast truncation algorithm for brain type conversions, we employ a third-party library [36] for half type conversions on MN4 exclusively because its CPUs do not provide native support. Note also that conversions made in software with this library inhibit the vectorization of some parts of our TVC library, notably while storing the results from vector–matrix multiplication kernels, vector updates (i.e., axpby functions), and global array reductions via MPI. As a result, the entire application performance plummets making it unusable. To mitigate this, we introduce an intermediate array to cache high precision results. The following code snippet corresponding to the axpby function illustrates this:
@@
Figure 7 shows the absolute performance of the dHOPM3 algorithm for various floating-point precision formats on two CPU architectures and up to eight processes. Starting with MN4, the throughput values corresponding to double precision are taken directly from Fig. 4(b). With single and single-double precision the throughput is practically increased by a factor of two regardless of the tensor order. Brain-single precision doubles the speed of single precision, yielding about 4×4$\times$4 × speedup with respect to double precision. The half-single precision results obtained with the half library are quite competitive with the previous case demostrating the advantages of caching. Interestingly, there is a considerable performance hit of $-34$% for $d=3$, which makes for an effective speedup of 3.1×3.1$\times$3.1 × over double precision. In general, the matrix case ($d=2$) yields slightly worse results for half-single and brain-single precision, which is attributed to the global array reduction involving tens of thousands of elements per process, all carried out in mixed precision arithmetic. By contrast, mixed precision dTVC does not exhibit such behaviour since it does not rely on collectives. Generally, dTVC and dHOPM3 mixed precision results show similar trends, but the latter constitutes a more challenging example due to its variety of kernels and synchronous collectives.
Figure 7(b) shows inconclusive results on the ARM architecture. On the one hand, using single precision does not necessarily double the performance as it was the case of MN4. On the other hand, and contrarily to what is observed in MN4, the single-double results show hindered performance for lower order tensors ($d<4$). Although the CPUs of CTE-ARM provide hardware support for half precision floating points, half-single results reflect a speedup below 2×2$\times$2 × w.r.t. single precision and about 3×3$\times$3 × w.r.t. double precision. Similarly to MN4, worse mixed precision speedup figures are associated with lower order tensors due to the increasing cost of mixed precision collectives. Finally, the brain-single results based upon a simple truncation made in software show a performance similar to half-single mixed precision arithmetic for most tensors ($d>3$).
Finally, we do not have mixed precision results on GPUs since, to the best of our knowledge, NCCL does not provide mixed precision MPI routines. Moreover, CUDA kernels do not offer the degree of flexiblity required to consistently integrate all the mixed precision kernels needed by the distributed higher order power method. In this regard, further work is needed to extend the support for mixed precision in HPC software, including message-passing interface (MPI) libraries.
@@ 6 Conclusions and future work
This work has presented a distributed tensor–vector contraction algorithm built on top of a native TVC shared-memory library recently published. It has also introduced a carefully optimized, distributed higher order power method, dHOPM3, which takes advantage of data-flow parallelization. The performance of dTVC is compromised when the splitting and contraction modes coincide and, while this can be easily circumvented for one dTVC over a particular contraction mode, it is unavoidable in the dHOPM algorithm where all modes are constricted. The analytical formulae derived from hypersquare tensors demonstrate that the best parallel performance is obtained when splittings break the contiguity of tensor elements, in contrast to a naive approach that seeks to preserve this property for simplicity.
Numerical experiments of dTVC and dHOPM3 algorithms have been carried out on three different architectures featuring CPUs, GPUs, and high end memory, using native and looped (cuBLAS-based) implementations. On CPUs, we obtain dTVC bandwidth figures about 50% to 80% of the theoretical peak using up to 128 MPI processes, which are on par with those reported by the popular STREAM benchmark on a single NUMA node. In the case of dHOPM3, figures are slightly lower, but overall they follow the same trend and, in absolute terms, the average throughput (kernel invocations per second) is, as expected, roughly half of that of dTVC regardless of the tensor order. On GPUs, the looped dTVC algorithm with state-of-the-art CUDA kernels brings an excess of variability into the results and, more importantly, evidences performance issues related to low device occupancy and increasing synchronous communication. In this regard, general-purpose computers are still able to compete against newer architectures equipped with low capacity HBM or GPUs on strong scalability scenarios. Lastly, it has been proved that the use of mixed precision kernels with our proposed caching technique is an effective way of almost doubling or even quadrupling the throughput of dTVC and dHOPM algorithms, although the real performance may remain architecture-dependent.
Future work will focus on the development of a CUDA-based, native TVC kernel for GPUs following the same philosophy adopted in our TVC library for CPUs. This will enable beyond state-of-the-art dTVC performance, and subsequently dHOPM performance, on these accelerators. On the other hand, it will be worth assessing the effects of multidimensional tensor splitting and the incursion of tensor unfolding operations. Analytical formulae for hypersquare tensors and numerical experiments with fine-grain parallelization will shed light in the pros and cons of increasing the number of splitting dimensions. Finally, we will seek to integrate this work within already existing frameworks, e.g. by bringing dense multilinear algebra support to ALP.

### v2

@@ 5.5 Mixed precision effects on performance
Since both dTVC and dHOPM remain memory bounded algorithms, mixed precision can be employed to further increase their throughput. In fact, mixed precision is a very popular technique actively used in deep learning applications [34, 35]. In this work, low precision is used exclusively as a storage format and we promote numbers to high precision, i.e. by doubling their bit representation, just before computing them. Therefore, every arithmetic operation, besides accumulations, is done in high precision. Calculated quantities are then converted back to their original storage precision. It is worth noting that mixed precision also affects communication and hence requires the development of ad-hoc MPI functions that enable high-precision scalar and array sums while keeping data transfers in low precision.
We consider three mixed precision formats: single-double, half-single, and brain-single. The half format stands for the IEEE 754-2008 binary16 standard while brain consists of a 16-bit truncation of the binary32 type from the same standard. While it is straightforward to create a fast truncation algorithm for brain type conversions, we employ a third-party library [36] for half type conversions on MN4 exclusively because its CPUs do not provide native hardware support. Note also that conversions made in software with this library inhibit the vectorization of some parts of our TVC library, notably while storing the results from vector–matrix multiplication kernels, vector updates (i.e., axpby functions), and global array reductions via MPI. As a result, the entire application performance plummets making it unusable. To mitigate this, we introduce an intermediate array to cache high-precision results. The following code snippet extracted from Ref. [10] and corresponding to the axpby function illustrates this:
@@
Figure 7 shows the absolute performance of the dHOPM3 algorithm for various floating-point precision formats on two CPU architectures and eight processes. Starting with MN4, the throughput values corresponding to double precision are taken directly from Fig. 5(b). With single and single-double precision the throughput is practically increased by a factor of two regardless of the tensor order. Brain-single precision doubles the speed of single precision, yielding about 4×4$\times$4 × speedup with respect to double precision. The half-single precision results obtained with the half library are quite competitive with the previous case demostrating the advantages of caching. Interestingly, there is a considerable performance hit of $-34$% for $d=3$, which makes for an effective speedup of 3.1×3.1$\times$3.1 × over double precision. In general, the matrix case ($d=2$) yields slightly worse results for half-single and brain-single precision, which is attributed to the global array reduction involving tens of thousands of elements per process, all carried out in mixed precision arithmetic. By contrast, mixed precision dTVC does not exhibit such behaviour since it does not rely on collectives. Generally, dTVC and dHOPM3 mixed precision results show similar trends, but the latter constitutes a more challenging example due to its variety of kernels and synchronous collectives.
Figure 7(b) shows inconclusive results on the ARM architecture. On the one hand, using single precision does not necessarily double the performance as it was the case of MN4. On the other hand, and contrarily to what is observed in MN4, the single-double results show hindered performance for lower-order tensors ($d<4$). Although the CPUs of CTE-ARM provide hardware support for half precision floating points, half-single results reflect a speedup below 2×2$\times$2 × w.r.t. single precision and about 3×3$\times$3 × w.r.t. double precision. Similarly to MN4, worse mixed precision speedup figures are associated with lower-order tensors due to the increasing cost of mixed precision collectives. Finally, the brain-single results based upon a simple truncation made in software show a performance similar to half-single mixed precision arithmetic for most tensors ($d>3$).
Finally, we do not have mixed precision results on GPUs since, to the best of our knowledge, NCCL does not provide mixed precision MPI routines. Moreover, CUDA kernels do not offer the degree of flexiblity required to consistently integrate all the mixed precision kernels needed by the distributed higher-order power method. In this regard, further work is needed to extend the support for mixed precision in HPC software, including message-passing interface (MPI) libraries.
@@ 6 Conclusions and future work
This work has presented a distributed tensor–vector contraction algorithm built on top of a CPU-native TVC shared-memory library recently published. It has also introduced a carefully optimized, distributed higher-order power method, dHOPM3, which takes advantage of data-flow parallelization. The performance of dTVC is compromised when the splitting and contraction modes coincide and, while this can be easily circumvented for one dTVC over a particular contraction mode, it is unavoidable in the dHOPM algorithm where all modes are constricted. The analytical formulae derived from hypersquare tensors demonstrate that the best parallel performance is obtained when the splitting breaks the contiguity of tensor elements, in contrast to a naive approach that seeks to preserve this property for simplicity.
Numerical experiments of dTVC and dHOPM3 algorithms have been carried out on three different architectures featuring CPUs, GPUs, and high-end memory, using native and looped (cuBLAS-based) implementations. On CPUs, we obtain dTVC bandwidth figures about 50% to 80% of the theoretical peak using up to 128 MPI processes, which are on par with those reported by the popular STREAM benchmark on a single NUMA node. In the case of dHOPM3, figures are slightly lower, but overall they follow the same trend and, in absolute terms, the average throughput (kernel invocations per second) is, as expected, roughly half of that of dTVC regardless of the tensor order. On GPUs, the looped dTVC algorithm with state-of-the-art CUDA kernels brings an excess of variability into the results and, more importantly, evidences performance issues related to low device occupancy and increasing synchronous communication. In this regard, general-purpose computers are still able to compete against newer architectures equipped with low-capacity HBM or GPUs on strong scalability scenarios. Lastly, it has been proved that the use of mixed precision kernels with our proposed caching technique is an effective way of almost doubling or even quadrupling the throughput of dTVC and dHOPM algorithms, although the real performance may remain architecture-dependent.
Future work will focus on the development of a CUDA-based, native TVC kernel for GPUs following the same philosophy adopted in our CPU library. This will enable beyond state-of-the-art dTVC performance, and subsequently dHOPM performance, on these accelerators. On the other hand, it will be worth assessing the effects of multidimensional tensor splitting and the incursion of tensor unfolding operations. Analytical formulae for hypersquare tensors and numerical experiments with fine-grain parallelization will shed light in the pros and cons of increasing the number of splitting dimensions. Finally, we will seek to integrate this work within already existing frameworks, e.g. by bringing dense multilinear algebra support to ALP.